write and solve an equation to find the measure of the third angle of a triangle with two angles measuring 32 degrees and 22 degrees. the sum of the angles of a triangle is 180 degrees.
Answer by funmath(2873)

write and solve an equation to find the measure of the third angle of a triangle with two angles measuring 32 degrees and 22 degrees. the sum of the angles of a triangle is 180 degrees.
Let the angle you don't know be: x
sum means +
x+32+22=180
x+54=180
x+54-54=180-54
x+0=126
x=126 degrees
:
Sanity check, if we add the two angles we were given with the angle that we got, does it = 180 degree?
126+32+22=180
180=180 Yes, we appear to be sane.
Happy Calculating!!!